OpenStudy (izzy18970):

Write an integer for each situation. A stock went up $4

OpenStudy (izzy18970):

A stock went up $4 Lex lost $5 out of his pocket An antique bowl gained $300 in value.

OpenStudy (izzy18970):

HELP please

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (izzy18970):

I finished that just now. But one question.

OpenStudy (izzy18970):



jimthompson5910 (jim_thompson5910):

what's your question with this one?

OpenStudy (izzy18970):

How much more did Trish make on her investment than Sammy?

jimthompson5910 (jim_thompson5910):

Trish made 24 dollars Sammy made -38 dollars (so she lost 38 dollars)

jimthompson5910 (jim_thompson5910):

subtract the two to get 24 - (-38) = 24 + 38 = 62 so Trish made $62 more than Sammy

OpenStudy (izzy18970):

So I did get it right!!!! In class I did that, but I thought I was wrong. i subtracted 24 - (-38)

jimthompson5910 (jim_thompson5910):

you did it right

OpenStudy (izzy18970):

And got 62

jimthompson5910 (jim_thompson5910):

62 is correct